# DataRater: Meta-Learned Dataset Curation

### v1

@@ 1 Introduction
Figure 1: Compute needed to achieve baseline performance using the DataRater. A baseline $1$B model is trained on the unfiltered dataset, while a second $1$B model is trained analogously on the same dataset, but filtered by a DataRater. The x-axis states the underlying dataset, with $3$ evaluation metrics each, while the y-axis shows the fraction of compute needed to match the baseline in grey. The overhead of using the DataRater for filtering online is shown in pink. ‘Validation set’ refers to the respective validation set of the underlying dataset. Net compute gain is shown in green. The figure shows that filtering data using the DataRater results in significant overall compute gains for lower quality datasets like the Pile & C4/noclean.
@@
  • DataRater framework. We introduce the DataRater, a novel meta-learning framework designed to estimate the value of individual data points for training foundation models. Our approach aims to automate dataset curation by meta-learning the value of data.
@@
  • Cross-scale generalisation of data valuation. Critically, a DataRater model meta-trained with fixed-scale inner models ($400$M parameters) effectively generalizes its learned data valuation policy to enhance training across a spectrum of target model scales ($50$M to $1$B parameters, also Figure 6). Moreover, optimal data discard proportions are shown to be consistent across these model scales (Figure 4).
  • Insight into data quality. Our analysis (Section  3.6) reveals that the DataRater learns to identify and down-weight data that aligns with human intuitions of low quality, such as incorrect text encodings, OCR errors, and irrelevant content.
@@ 2 Meta-Learned Dataset Curation
Figure 2: DataRater meta-learning schematic. This diagram shows how the DataRater is updated using meta-gradients to minimise an outer loss, by back-propagating through multiple inner model updates (we experimented with up to $8$ inner updates) over weighted inner data batches. For more details on the implementation see Algorithm 1.
@@
3:$\eta_{0}\leftarrow\textsc{InitMetaparams}()$triangleright Initialise DataRater.
@@
10:         $\mathcal{B}_{k}\sim\mathcal{P}_{N_{\text{outer}}}(\mathcal{D}_{\text{test}})$triangleright Sample outer batch.
11:         $g_{k}^{i}=\frac{1}{N_{\text{outer}}}\sum_{\mathbf{x}\in\mathcal{B}_{k}}\nabla_{\eta}L(\mathbf{x};\theta_{k+1}^{i}(\eta_{k}))$triangleright Form meta-gradient.
12:         $\bar{\eta}^{i}=H(\eta_{k},g_{k}^{i})$triangleright Compute per-model meta-parameter update.
13:     $\eta_{k+1}=\frac{1}{N_{\text{models}}}\sum_{i}\bar{\eta}^{i}$triangleright Update DataRater.
14:Return $\eta_{K}$triangleright Optimised DataRater meta-parameters.
@@
18:         $\mathcal{B}_{t}\sim\mathcal{P}_{N_{\text{inner}}}(\mathcal{D}_{\text{train}})$triangleright Sample inner batch.
19:         $g_{t}=\sum_{\mathbf{x}\in\mathcal{B}_{t}}\nabla_{\theta}\ell(\mathbf{x};\theta_{t})\cdot\sigma_{\mathcal{B}_{t}}(\phi_{\eta}(\mathbf{x}))$triangleright Form weighted gradient using data ratings.
20:         $\theta_{t+1}=\Theta(\theta_{t},g_{t})$triangleright Update parameters.
21:     Return $\theta_{T}$triangleright $T$-steps updated inner model parameters.
@@
Meta-optimisation. To optimise the parameters, $\eta$, of a DataRater model approximately, we use a meta-gradient method (Xu et al., 2018): $\eta$ is optimised via a stochastic gradient method, where the gradient of the loss $L$ (which is called outer loss in this context) with respect to $\eta$ is computed by back-propagating through multiple optimiser updates for $\theta$ with respect to the so-called inner loss $\ell$, as illustrated in Figure 2; the method is given formally in Algorithm $1$Pseudo-code omits optimiser states and periodic resetting of inner model parameters..
@@
We instantiate the DataRater as a non-causal transformer and optimise its parameters $\eta$ using meta-gradient descent; i.e. back-propagating through the unrolled optimisation of the inner model. We back-propagate through a truncated window of $2$ inner model updates to compute the meta-gradient [^2] We experimented with $1$, $2$, $4$ and $8$ inner model updates, and found that using $2$ inner updates results in very similar DataRater performance as using more.. This requires computing vector products with second-order derivatives (Hessian matrices), which can be very challenging computationally. To make this computation feasible, we use the scalable bilevel optimisation reparameterisation of Kemaev et al. (2025) titled MixFlow-MG. This method exploits symmetries present in bilevel gradient optimisation problems, using mixed-mode differentiation to significantly reduce RAM (HBM) usage. It enables us to optimise DataRater meta-parameters efficiently – even for $50$M DataRater models and $400$M inner models, while back-propagating through several inner model update steps.
To stabilise meta-gradient computation we: (1) use a population of inner models; (2) pass the meta-gradient for each inner model through an individual Adam optimiser (Kingma and Ba, 2015), averaging the resulting meta-gradient updates; (3) periodically reinitialise inner models to stratify learning progress.
@@
For data curation, with a trained DataRater, we use top-$K$ filtering at the batch-level: instead of weighting data, we remove data with low predicted value. To implement top-K filtering at the batch level, for a given target discard fraction $\rho\in[0,1)$, and a nominal batch size $N$, we upsample a larger batch of data points sized $N\frac{\rho}{1-\rho}$, score them using a DataRater, and filter out the bottom $(100\cdot\rho)$-percent of samples.
DataRater scores can also be equivalently used to select data points individually – without requiring access to a batch of data. This requires access to the CDF of the ratings distribution, $F_{\Phi}:\mathbb{R}\to[0,1]$. To emulate the batch-level decision for top-$K$ filtering given a batch size of $B$, for a data point $\mathbf{x}$ with $p=F_{\Phi}(\phi(\mathbf{x}))$, we can calculate the probability of keeping it as $\text{P}_{\text{accept}}(\mathbf{x})=\sum_{s=0}^{K-1}(1-p)^{s}p^{B-s-1}$ (since we keep a data point if select at most $K-1$ “better” points in a batch). This is useful for building massively parallel data filtering pipelines using, e.g., Apache Beam (Dean and Ghemawat, 2004), as filtering decisions can be made locally.
@@ 3 Experimental Results
To evaluate the effectiveness of our method in curating a dataset we perform the following three main steps: (1) meta-learn a DataRater for the input dataset $\mathcal{D}$; (2) use the trained DataRater to curate it, forming $\mathcal{D}^{*}$; (3) train language models (of $50$M, $150$M, $400$M and $1$B sizes) from random initialisation on the curated dataset and subsequently evaluate their performance.
@@
DataRater configuration. For each of the three datasets, we train a distinct $50$M parameter non-causal transformer as the DataRater model. Each DataRater is meta-learned using a population of eight $400$M parameter inner language models (see Appendix A.1 for meta-learning checkpoint selection details). Importantly, for each input dataset, the inner models and their corresponding DataRater model are optimised on disjoint subsets of that dataset’s training data.
@@
We measure negative log likelihood (NLL) on the validation set of each of the three input datasets, as well as on (English-language) Wikipedia. We measure accuracy on the following downstream tasks: HellaSwag (Zellers et al., 2019), SIQA (Sap et al., 2019), PIQA (Bisk et al., 2020), ARC Easy (Clark et al., 2018) and Commonsense QA (Talmor et al., 2019), which we selected as they provide useful signal at the model scales we use (i.e. benchmark performance increases during the course of training and across model scales). In Figure 1 we plot the fraction of compute needed to match baseline performance for $1$B models; this is calculated in terms of the number of training steps on the DataRater-filtered dataset needed to reach the final baseline performance. We leave this metric undefined for cases where training on the filtered datasets does not reach baseline performance (e.g. for some cases on the C4 dataset).
Figure 5: Learning curves of 1B models trained on three separate underlying datasets, with DataRater filtering and without. On both C4/noclean and the Pile, models trained on DataRater filtered data match the baseline’s final performance while using considerably fewer training steps, while also resulting in improved final performance. Using a DataRater to filter C4, which is a considerably higher quality dataset, results in minor performance improvements on average (i.e. see Table 1 for full results) while exhibiting trade-offs in downstream evaluations as exemplified here.
@@ 3.3 Can the DataRater accelerate learning?
The DataRater indeed accelerates learning, resulting in considerable FLOPS savings for lower quality datasets like the Pile and C4/noclean. For $1$B models, Figure 1 shows precisely the amount of compute saved by using the DataRater, when compared to baseline training, while also showing the compute overhead due to DataRater filtering. Figure 5 shows performance metrics during training of $1$B models. This shows that using DataRater filtered datasets results in improved final performance, and not just in faster learning.
Of course, training the DataRater model itself must also be accounted for. Meta-training a DataRater requires approximately $58.4\%$ of the FLOPS needed to train a single $1$B LLM. Given that the filtered dataset produced by the DataRater will be used to train numerous other LLMs, of possibly much larger scale, we argue that the cost of training the DataRater can be amortised away.
@@ 3.4 How much data should be removed?
To select the best discard proportion for each dataset, we run a sweep over $5$ candidate discard proportions ($10\%$, $25\%$, $50\%$, $75\%$ and $90\%$), using the smallest model ($50$M) and select the discard proportion resulting in the best validation set NLL. In Figure 4 we show that using the smallest model size is sufficient, as the best discard proportion is shared across model sizes. The optimal value depends on the underlying data quality: for C4 we drop $10\%$, for C4/noclean $50\%$ and for the Pile $75\%$.
@@ 3.5 Is the DataRater robust to model scale?
We trained a single DataRater per dataset, with a fixed inner model size of $400$M. The DataRater generalises across model scales: in experiments over $3$ input datasets over $4$ model scales and $7$ metrics, $73/84$ showed improvements (see Figure 6). For the lower quality datasets, C4/noclean and the Pile, the performance improvements are consistent over model scale for NLL metrics and HellaSwag, whereas downstream evaluations have more variance.
@@ 4 Related Work
Bilevel Optimisation and Meta-Learning for Data Selection. The most analogous approaches to the DataRater frame data selection as a bilevel optimisation problem to learn a selection or weighting policy. Maclaurin et al. (2015) demonstrated computing exact gradients of validation performance with respect to hyperparameters by reversing learning dynamics. Many works (e.g., (Pedregosa, 2016; Lorraine et al., 2020)) make use of the implicit function theorem to efficiently approximate the solution to similar bilevel optimisation problems. Differentiable Data Selection (Wang et al., 2020) explicitly uses a bilevel setup where a scorer network learns to weight data instances to optimise a main model’s performance on a development set, using gradient alignment as a reward. Other works explicitly learn training data distributions under domain shift using similar bilevel frameworks (Grangier et al., 2024).
Contemporaneously, SEAL (Shen et al., 2025) learns a ‘data ranker’ via bilevel optimisation to select fine-tuning data to enhance LLM safety while preserving utility; our formulation allows for arbitrary inner and outer losses, but we focus empirical evaluation on improving training efficiency. SEAL optimises the resulting bilevel optimisation problem using a penalty method (Clarke, 1990), while we use exact meta-gradients. Concurrent work (Engstrom et al., 2025) similarly applies meta-gradients to optimise training configurations, including to select multi-modal data for CLIP (Radford et al., 2021) and instruction tuning data for Gemma-2B (Gemma Team, 2024) with a similar meta-objective as our method. Their method tracks one meta-parameter per data point, while we use function approximation to learn to assign value to arbitrary individual data points. We use all techniques from MixFlow-MG (Kemaev et al., 2025) (gradient check-pointing, mixed-differentiation, etc.) to efficiently compute meta-gradients, while their method introduces the ‘REPLAY’ algorithm for the same purpose.

### v2

@@ 1 Introduction
Figure 1: Compute needed to achieve baseline performance using the DataRater. A baseline 11B model is trained on the unfiltered dataset, while a second 11B model is trained analogously on the same dataset, but filtered by a DataRater. The x-axis states the underlying dataset, with 33 evaluation metrics each, while the y-axis shows the fraction of compute needed to match the baseline in grey. The overhead of using the DataRater for filtering online is shown in pink. To calculate this metric we convert both the LLM training step cost and the DR inference step costs into FLOPs (while also accounting for batch-size oversampling). ‘Validation set’ refers to the respective validation set of the underlying dataset. Net compute gain is shown in green. The figure shows that filtering data using the DataRater results in significant overall compute gains for lower quality datasets like the Pile & C4/noclean.
@@
  • DataRater framework. We introduce the DataRater, a novel meta-learning framework designed to estimate the value of individual data points for training foundation models, relative to a specified meta-objective. Our approach aims to automate dataset curation by meta-learning the value of data.
@@
  • Cross-scale generalisation of data valuation. Critically, a DataRater model meta-trained with fixed-scale inner models (400400M parameters) effectively generalizes its learned data valuation policy to enhance training across a spectrum of target model scales (5050M to 11B parameters, also Figure 6). Moreover, optimal data discard proportions are shown to be consistent across these model scales (Figure 4).
  • Insight into data quality. Our analysis (Section 3.6) reveals that the DataRater learns to identify and down-weight data that aligns with human intuitions of low quality, such as incorrect text encodings, OCR errors, and irrelevant content.
@@ 2 Meta-Learned Dataset Curation
Figure 2: DataRater meta-learning schematic. This diagram shows how the DataRater is updated using meta-gradients to minimise an outer loss, by back-propagating through multiple inner model updates (we experimented with up to 88 inner updates) over weighted inner data batches. For more details on the implementation see Algorithm 1.
@@
[…]
@@
10:   $\mathcal{B}_{k}\sim\mathcal{P}_{N_{\text{outer}}}(\mathcal{D}_{\text{test}})$ $\triangleright$ Sample outer batch.
11:   $g_{k}^{i}=\frac{1}{N_{\text{outer}}}\sum_{\mathbf{x}\in\mathcal{B}_{k}}\nabla_{\eta}L(\mathbf{x};\theta_{k+1}^{i}(\eta_{k}))$ $\triangleright$ Form meta-gradient.
12:   $\bar{\eta}^{i}=H(\eta_{k},g_{k}^{i})$ $\triangleright$ Compute per-model meta-parameter update.
13:  $\eta_{k+1}=\frac{1}{N_{\text{models}}}\sum_{i}\bar{\eta}^{i}$ $\triangleright$ Update DataRater.
14:Return $\eta_{K}$ $\triangleright$ Optimised DataRater meta-parameters.
@@
18:   $\mathcal{B}_{t}\sim\mathcal{P}_{N_{\text{inner}}}(\mathcal{D}_{\text{train}})$ $\triangleright$ Sample inner batch.
19:   $g_{t}=\sum_{\mathbf{x}\in\mathcal{B}_{t}}\nabla_{\theta}\ell(\mathbf{x};\theta_{t})\cdot\sigma_{\mathcal{B}_{t}}(\phi_{\eta}(\mathbf{x}))$ $\triangleright$ Form weighted gradient using data ratings.
20:   $\theta_{t+1}=\Theta(\theta_{t},g_{t})$ $\triangleright$ Update parameters.
21:  Return $\theta_{T}$ $\triangleright$ $T$-steps updated inner model parameters.
@@
Meta-optimisation. To optimise the parameters, $\eta$, of a DataRater model approximately, we use a meta-gradient method (Xu et al., 2018): $\eta$ is optimised via a stochastic gradient method, where the gradient of the loss $L$ (which is called outer loss in this context) with respect to $\eta$ is computed by back-propagating through multiple optimiser updates for $\theta$ with respect to the so-called inner loss $\ell$, as illustrated in Figure 2; the method is given formally in Algorithm 1111Pseudo-code omits optimiser states and periodic resetting of inner model parameters..
@@
We instantiate the DataRater as a non-causal transformer and optimise its parameters $\eta$ using meta-gradient descent; i.e. back-propagating through the unrolled optimisation of the inner model. We back-propagate through a truncated window of $2$ inner model updates to compute the meta-gradient [^2] We experimented with $1$, 22, 44 and $8$ inner model updates, and found that using $2$ inner updates results in very similar DataRater performance as using more.. This process, handled implicitly by automatic differentiation, involves operations with second-order derivatives, such as Hessian matrices, which in practice are computationally expensive. To make this computation feasible, we use the scalable bilevel optimisation reparameterisation of Kemaev et al. (2025) called MixFlow-MG. This method exploits implicit symmetries present in bilevel gradient optimisation problems and uses mixed-mode differentiation to significantly reduce RAM (HBM) usage. In particular, we use all the core techniques from MixFlow-MG, including block-level rematerialisation with saving inner-level gradients, and mixed-mode differentiation itself. This enables us to optimise DataRater meta-parameters efficiently – even for 5050M DataRater models and 400400M inner models, while back-propagating through multiple inner model update steps.
To stabilise meta-gradient computation we: (1) use a population of inner models; (2) pass the meta-gradient for each inner model through an individual Adam optimiser (Kingma and Ba, 2015), averaging the resulting meta-gradient updates; (3) periodically reinitialise inner models to stratify learning progress. In the appendix in Figure 12 we plot the norm of the meta-gradient updates over the course of meta-training, for each inner model, showing that they remain stable and bounded. Furthermore, in Figure 13 we show that the temporal auto-correlation of DataRater scores approaches >0.95>0.95 after a few thousand meta-update steps, indicating that meta-training is converging.
@@
For data curation, with a trained DataRater, we use top-$K$ filtering at the batch-level: instead of weighting data, we remove data with low predicted value. To implement top-K filtering at the batch level, for a given target discard fraction $\rho\in[0,1)$, and a nominal batch size $N$, we upsample a larger batch of data points sized $\frac{N}{1-\rho}$, score them using a DataRater, and filter out the bottom $(100\cdot\rho)$-percent of samples.
DataRater scores can also be equivalently used to select data points individually – without requiring access to a batch of data. This requires access to the CDF of the ratings distribution, $F_{\Phi}:\mathbb{R}\to[0,1]$. To emulate the batch-level decision for top-$K$ filtering given a batch size of $B$, for a data point $\mathbf{x}$ with $p=F_{\Phi}(\phi(\mathbf{x}))$, we can calculate the probability of keeping it as $\text{P}_{\text{accept}}(\mathbf{x})=\sum_{s=0}^{K-1}\binom{B-1}{s}(1-p)^{s}p^{B-s-1}$ (since we keep a data point if select at most K1K-1 “better” points in a batch). This is useful for building massively parallel data filtering pipelines using, e.g., Apache Beam (Dean and Ghemawat, 2004), as filtering decisions can be made locally.
@@ 3 Experimental Results
To evaluate the effectiveness of our method in curating a dataset we perform the following three main steps: (1) meta-learn a DataRater for the input dataset $\mathcal{D}$; (2) use the trained DataRater to curate it, forming $\mathcal{D}^{*}$; (3) train language models (of 5050M, 150150M, 400400M and 11B sizes) from random initialisation on the curated dataset and subsequently evaluate their performance.
@@
DataRater configuration. For each of the three datasets, we train a distinct 5050M parameter non-causal transformer as the DataRater model. Each DataRater is meta-learned using a population of eight 400400M parameter inner language models (see Appendix A.1 for meta-learning checkpoint selection details). Importantly, for each input dataset, the inner models and their corresponding DataRater model are optimised on disjoint subsets of that dataset’s training data.
@@
We measure negative log likelihood (NLL) on the validation set of each of the three input datasets, as well as on (English-language) Wikipedia. We measure accuracy on the following downstream tasks: HellaSwag (Zellers et al., 2019), SIQA (Sap et al., 2019), PIQA (Bisk et al., 2020), ARC Easy (Clark et al., 2018) and Commonsense QA (Talmor et al., 2019), which we selected as they provide useful signal at the model scales we use (i.e. benchmark performance increases during the course of training and across model scales). In Figure 1 we plot the fraction of compute needed to match baseline performance for 11B models. To calculate this metric we convert both the LLM training step cost into FLOPs as well as the DR inference step cost into FLOPs, while also accounting for batch-size oversampling prior to DR online filtering. We leave this metric undefined for cases where training on the filtered datasets does not reach baseline performance (e.g. for some cases on the C4 dataset).
@@ 3.3 Can the DataRater accelerate learning?
The DataRater indeed accelerates learning, resulting in considerable FLOPS savings for lower quality datasets like the Pile and C4/noclean. For 11B models, Figure 1 shows precisely the amount of compute saved by using the DataRater, when compared to baseline training, while also showing the compute overhead due to DataRater filtering. Figure 5 shows performance metrics during training of 11B models. This shows that, particularly for lower-quality datasets, using DataRater filtered datasets often results in improved final performance, and not just in faster learning.
Of course, training the DataRater model itself must also be accounted for. Meta-training a DataRater requires approximately $58.4\%$ of the FLOPS needed to train a single 11B LLM. Given that the filtered dataset produced by the DataRater will be used to train numerous other LLMs, of possibly much larger scale, we argue that the cost of training the DataRater can be amortised away.
In Section A.2 in the appendix we compare the DataRater approach with a perplexity-based filtering method adapted from prior work (Ankner et al., 2025), showing that the DataRater results in better performance in a large majority of evaluations, i.e. in $16$ out of $21$ cases.
Figure 5: Learning curves of 1B models trained on three separate underlying datasets, with DataRater filtering and without. On both C4/noclean and the Pile, models trained on DataRater filtered data match the baseline’s final performance while using considerably fewer training steps, while also resulting in improved final performance. Using a DataRater to filter C4, which is a considerably higher quality dataset, results in minor performance improvements on average (i.e. see Table 2 in appendix for full results) while exhibiting trade-offs in downstream evaluations as exemplified here.
@@ 3.4 How much data should be removed?
To select the best discard proportion for each dataset, we run a sweep over $5$ candidate discard proportions ($10\%$, $25\%$, $50\%$, $75\%$ and $90\%$), using the smallest model (5050M) and select the discard proportion resulting in the best validation set NLL. In Figure 4 we show that using the smallest model size is sufficient, as the best discard proportion is shared across model sizes. The optimal value depends on the underlying data quality: for C4 we drop $10\%$, for C4/noclean $50\%$ and for the Pile $75\%$.
@@ 3.5 Is the DataRater robust to model scale?
We trained a single DataRater per dataset, with a fixed inner model size of 400400M. The DataRater generalises across model scales: in experiments over $3$ input datasets over $4$ model scales and $7$ metrics, $73/84$ showed improvements (see Figure 6). For the lower quality datasets, C4/noclean and the Pile, the performance improvements are consistent over model scale for NLL metrics and HellaSwag, whereas downstream evaluations have more variance.
@@ 3.6 What does the DataRater learn to do?
In Section A.4 in the appendix we perform a correlation analysis between the DataRater score and common language filtering heuristics, with results shown in Figure 14. The DataRater score is most positively correlated with the number of packed sub-sequences as well as with various metrics of text length; it is most negatively correlated with the proportion of non-alphanumeric characters, of punctuation and of unique characters. An OLS model results in an $R^{2}$ of $0.766$. As many of the heuristics are collinear, we fit a cross-validated Lasso ($L1$) model on Z-scored features to identify a small but predictive subset of heuristics. The Lasso model explains $75.3\%$ of the variance in the DataRater scores while using only $11$ non-zero coefficients. We find that the DataRater score can mostly be explained by the number of packed sub-sequences, the proportion of non-alphanumeric characters, the word count, the sequence length, the packing density as well as a few other heuristics. Since neither regression model can perfectly predict the DataRater’s score, it is likely the DataRater is identifying valuable patterns in the data that are not captured by these simple heuristics.
@@ 4 Related Work
Bilevel Optimisation and Meta-Learning for Data Selection. The most analogous approaches to the DataRater frame data selection as a bilevel optimisation problem to learn a selection or weighting policy. Maclaurin et al. (2015) demonstrate computing exact gradients of validation performance by reversing learning dynamics, while many subsequent works (e.g., (Pedregosa, 2016; Lorraine et al., 2020)) use the implicit function theorem for efficient approximation. Differentiable Data Selection (Wang et al., 2020) uses a bilevel setup where a scorer network learns to weight data instances to optimise a model’s performance, using gradient alignment as a reward. A similar bilevel weighting approach was used by (Guo et al., 2020) for safe semi-supervised learning.Others use similar bilevel frameworks to learn training data distributions under domain shift (Grangier et al., 2024) or to adaptively re-weight per-task gradients to improve multilingual training (Li and Gong, 2021).
Contemporaneously, SEAL (Shen et al., 2025) learns a ’data ranker’ via bilevel optimisation to select fine-tuning data to enhance LLM safety while preserving utility; our formulation mathematically allows for arbitrary inner and outer losses, but, in this work, we focus on improving training efficiency specifically. SEAL optimises the bilevel problem using a penalty method (Clarke, 1990), while we use exact meta-gradients. Concurrent work (Engstrom et al., 2025) similarly applies meta-gradients to optimise training configurations, including to select multi-modal data for CLIP (Radford et al., 2021) and instruction tuning data for Gemma-2B (Gemma Team, 2024) with a similar meta-objective as our method. Their method tracks one meta-parameter per data point, while we use function approximation to learn to assign value to arbitrary individual data points. We use all techniques from MixFlow-MG (Kemaev et al., 2025) (gradient checkpointing, mixed-differentiation, etc.) to efficiently compute meta-gradients, while their method introduces the ‘REPLAY’ algorithm for the same purpose.